OpenStudy (dmezzullo):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432

OpenStudy (anonymous):

All you have to do is to isolate your x....

OpenStudy (dmezzullo):

?

OpenStudy (anonymous):

What you do to one side, you do to the other! SO! x-0.8x = 0.8x-0.8x+72 .2x = 7x Find x.

OpenStudy (anonymous):

ahh!! .2x = 72***************** My bad

OpenStudy (dmezzullo):

so freakin lost

OpenStudy (dmezzullo):

is it a?

OpenStudy (anonymous):

288

OpenStudy (anonymous):

ITS THE SECOND ONE

OpenStudy (anonymous):

You have x's on both sides. What you want, is to have x's ONLY on one side! There is a rule that you can do anything to both sides, it will still be the same. FOR EXAMPLE: 4=4 divide bth sides by 4 1=1 OR 3=3 3-1=3-1 2=2 THIS SAME RULE APPLIES HERE

OpenStudy (anonymous):

yay another medal lol

OpenStudy (anonymous):

No, it's not a

OpenStudy (anonymous):

It's not b....

OpenStudy (dmezzullo):

who's right?

OpenStudy (anonymous):

It's c)

OpenStudy (anonymous):

72/.2 = 360

OpenStudy (dmezzullo):

u sure?

OpenStudy (anonymous):

I'm not 99% sure, I'm 110% sure. I've done this for years!

OpenStudy (anonymous):

You substract 0.8x from both sides so then it becomes 0.2x = 72 Then divide 72 by 0.2 72/0.2 = 360

OpenStudy (dmezzullo):

ty

OpenStudy (anonymous):

wowo its B

OpenStudy (anonymous):

And how did you get b?

OpenStudy (anonymous):

What I thought.

OpenStudy (anonymous):

Lol, wow.

OpenStudy (anonymous):

If I could, I'd give ya all a medal.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

You can, you give it , tanke it back, give it to someone else, etc.